OpenStudy (anonymous):

Solve and graph the absolute value inequality: |4x + 1| ≤ 5

OpenStudy (anonymous):

start with \[-5\leq 4x+1\leq 5\] and solve in two steps 1) subtract 1 2) divide by 4

OpenStudy (anonymous):

solve first, graph second

OpenStudy (anonymous):

there are 4 graphs

OpenStudy (anonymous):

you won't know what the graph is supposed to look like until you solve it

OpenStudy (anonymous):

yeah, i know, it is multiple choice what i am saying is, you need to solve it first, because otherwise you will not know what the graph is supposed to look like

OpenStudy (anonymous):

@AllTheLonelyKillers have you done these before like with graphs?

OpenStudy (anonymous):

alright :]

OpenStudy (anonymous):

i think so maybe

hero (hero):

@MyChem...@satellite73 knows what he's talking about. Listen to HIM

OpenStudy (anonymous):

\[-5\leq 4x+1\leq 5\] subtract 1\[-6\leq 4x\leq 4\] divide by \(4\)\ \[-\frac{6}{4}\leq x\leq 1\] or \[-1.5\leq x\leq 1\]

OpenStudy (anonymous):

now that we know what the inequality is, we know what the graph looks like the stuff shaded between \(-1.5\) and \(1\) with closed circles at each end

OpenStudy (anonymous):

you can't solve for the graph without first solving the inequality in other words, you cannot look at the graph and say "oh, it is this one" without first solving for \(x\)

OpenStudy (anonymous):

alright :]

OpenStudy (anonymous):

http://www.youtube.com/watch?v=HgWjFPQYjBE

OpenStudy (anonymous):

@Hero thanks for the vote of confidence! appreciated

OpenStudy (anonymous):

O.O

OpenStudy (anonymous):

we are now best friends! <3

OpenStudy (anonymous):

of course

OpenStudy (anonymous):

I love gerard way! :D